# Survey of Video Diffusion Models: Foundations, Implementations, and Applications

### v1

@@ 4.5 Long video
The world exists in three spatial dimensions plus time, while a video represents a 2D projection of this 3D world. Consequently, video diffusion models should adhere to certain 3D constraints. These models generate multiple frames simultaneously and capture inter-frame relationships through attention layers spanning different frames. Existing research, such as SORA (Li et al., 2024h), demonstrates that by training on large-scale datasets, networks can implicitly learn 3D priors embedded within the data.
@@ 4.6.1 Training on 3D dataset
SV3D (Blattmann et al., 2023a) fine-tunes a pre-trained stable video diffusion model on orbiting videos of 3D synthetic objects from the GSO (Downs et al., 2022) and OmniObject3D (Wu et al., 2023d) datasets. These orbiting videos include both static orbits, where the camera circles an object at regularly spaced azimuths with a fixed elevation, and dynamic orbits, where the camera trajectory is irregular. Similarly, V3D (Chen et al., 2024f) fine-tunes on the Objaverse dataset, leveraging 360-degree orbiting videos, along with proprietary datasets featuring similar setups (Junlin Han, 2024; Melas-Kyriazi et al., 2024).
@@
CineMaster (Wang et al., 2025) proposes a customized data labeling pipeline that leverages existing depth prediction (Yang et al., 2024e; f), instance segmentation (Ravi et al., 2024) and scene reconstruction (Zhang et al., 2024d) models to generate labeled data. This enables a 3D-aware model to train on large-scale, in-the-wild video data.
@@ 4.6.2 Architecture for 3D diffusion models
Recently, multi-view diffusion models (Shi et al., 2023d; Tang et al., 2023; Yang et al., 2024b; Liu et al., 2024b; f; Hu et al., 2024; Kant et al., 2024; Yang et al., 2024d; Oh et al., 2023; Tang et al., 2024) have gained significant attention for static 3D scene generation. By employing cross-frame attention, these models effectively enhance the spatial relationships necessary for 3D consistency. For static 3D scene generation, epipolar-constrained attention (Xu et al., 2024b; Huang et al., 2023c; Zheng et al., 2024a) further enforces strict adherence to 3D constraints, ensuring greater structural accuracy.
@@
In DiT-like structures (Gao*et al., 2024; Chou et al., 2024; Liang et al., 2024b; Gu et al., 2025; Wang et al., 2025), cross-frame attention is inherently achieved as tokens for different temporal and spatial locations are jointly mixed through transformer, allowing the model to learn 3D priors directly from data.
@@ 4.6.3 Camera conditioning
  • MLP-based embedding. A straightforward method is to use a fully connected multi-layer perceptron (MLP) to embed camera parameters. These parameters are often represented as elevation and azimuth angles, as in SV3D (Voleti et al., 2024; Yang et al., 2024c), panning and zooming movement (Yang et al., 2024h), or as a relative camera extrinsic matrix (Van Hoorick et al., 2024; Wang et al., 2024k; 2025). This embedding serves as a global condition, typically introduced into the diffusion model via cross-attention layers.
@@ 5.1 Video representation learning
As shown by previous works in image generation (Chi et al., 2024), a pretrained video generative model might be able to serve as prior, loss, or regularize in learning video representations. Recently, a pioneer work (Yeh et al., 2024) employs image and video generative models for hateful visual content recognition. It queries well-trained SOTA image and video generative models to automatically scale up training data pairs for hateful visual content recognition, achieving superior performance.
@@ 5.2 Video retrieval
Video retrieval, mainly video-text retrieval (VTR), which involves cross-modal alignment and abstract understanding of temporal images (videos), has been a popular and fundamental task of language-grounding problems (Wang et al., 2020a; b; 2021; Yu et al., 2023b). Most existing conventional video-text retrieval frameworks (Yu et al., 2017; Dong et al., 2019; Zhu & Yang, 2020; Miech et al., 2020; Gabeur et al., 2020; Dzabraev et al., 2021; Croitoru et al., 2021) focus on learning powerful representations for video and text and extracting modality-dependent representations with modality-dependent encoders. Inspired by the success of self-supervised pretraining methods (Devlin et al., 2019; Radford et al., 2019; Brown et al., 2020) and vision-language pretraining (Li et al., 2020; Gan et al., 2020; Singh et al., 2022) on large-scale unlabeled cross-modal data, recent works (Lei et al., 2021; Cheng et al., 2021; Gao et al., 2021; Ma et al., 2022; Park et al., 2022; Wang et al., 2022a; b; Zhao et al., 2022; Gorti et al., 2022; Luo et al., 2022; Liu et al., 2022d; a; Cao et al., 2022) have attempted to pretrain or fine-tune video-text retrieval models (Gorti et al., 2022; Lei et al., 2021).
With the rapid progress in video generation, leveraging the advancement in video generation (Chen et al., 2024b; Muaz et al., 2023) to advance video retrieval by reranking (Zhou et al., 2024e; Dutta et al., 2021) or other methods becomes more and more promising. On the other side, retrieval augmented generation could also help faithful video generation (Arefeen et al., 2024; Rosa, 2024; Tevissen et al., 2024; He et al., 2023a).
@@ 5.4.1 Video diffusion for 3D generation
V3D (Chen et al., 2024f) leverages the predictions of fine-tuned video diffusion models as direct supervision to generate 3D Gaussian splatting or meshes. SV3D (Voleti et al., 2024) employs the outputs of a fine-tuned Stable Video Diffusion model to supervise a Neural Radiance Field (NeRF) in the coarse stage, then uses SDS loss to refine high-frequency details of DMTet (Shen et al., 2021) initialized from the coarse stage. IM-3D (Melas-Kyriazi et al., 2024) adopts an instructive NeRF-to-NeRF (Haque et al., 2023) pipeline for distillation. This involves iteratively adding noise to rendered results and using a diffusion model to denoise the images, which then serve as targets for updating the 3D representation. VFusion3D (Junlin Han, 2024) demonstrates the potential of fine-tuned video diffusion models as data generators for creating large-scale 3D datasets. These datasets can be used to train feed-forward 3D generation models. Hi3D (Yang et al., 2024c) follows a coarse-to-fine 3D generation pipeline. It first generates low-resolution orbiting videos using a text-to-video model, then refines these into high-resolution orbiting videos with a video-to-video model. Both models are fine-tuned from pre-trained video diffusion models with camera condition extensions. The high-resolution orbiting videos are subsequently used to reconstruct 3D meshes. ReconX (Liu et al., 2024a) fine-tunes video diffusion models for sparse-view novel view synthesis, enabling efficient reconstruction of 3D scenes from limited input views. With recent advancements in large reconstruction models (LRMs) (Hong et al., 2023b; Zhang et al., 2024e), Wonderland (Liang et al., 2024b) demonstrates that by training a feed-forward 3D reconstruction model from the latent space of a video generative model, remarkable scene generation results can be achieved. Some examples are shown in Fig. 20.
@@ 5.4.2 Video diffusion for 4D generation
Several approaches leverage video diffusion models to train dynamic Neural Radiance Fields (NeRFs) or other 4D representations: MAV3D (Singer et al., 2023), 4D-fy (Bahmani et al., 2024c), Dream-in-4D (Zheng et al., 2024b), Align-Your-Gaussians (Ling et al., 2024a), and Animate124 (Zhao et al., 2023e) apply Score Distillation Sampling (SDS) loss along the temporal axis to train dynamic NeRFs using video diffusion models. These methods enable the models to learn motion dynamics directly from video diffusion. Vidu4D (Wang et al., 2024g) and STAG4D (Zeng et al., 2024b) propose robust reconstruction pipelines to reconstruct dynamic Gaussian Surfels or Gaussian splatting with monocular videos generated by video diffusion models. PLA4D (Miao et al., 2024) animates static 3D Gaussian splatting generated by 3D diffusion models, using video diffusion predictions combined with a motion alignment model. TC4D (Bahmani et al., 2024a) uses SDS loss to supervise dynamic NeRFs, incorporating a trajectory-conditioned video model that disentangles global and local motion components, facilitating the generation of large-scale motions.
Dynamic motion knowledge can also be combined with spatial 3D knowledge from multi-view diffusion models: EG4D (Sun et al., 2024) use a combination of SVD and SV3D models to directly generate multi-view videos, which are then used to supervise a 4D Gaussian splatting. 4Real (Yu et al., 2024a) generates 4D videos by using video diffusion model to supervise fixed-viewpoint video and freeze-time video with SDS loss. The use of deformable 3D GS ensures the temporal and view consistency. DreamGaussian4D (Ren et al., 2023) employs a training-free video-to-video pipeline that uses pre-trained video diffusion models to refine the dynamic textures of generated mesh sequences.
@@ 6 Conclusion and outlook
Corresponding to these challenges, we anticipate several promising research directions: 1) the development of more efficient architectures and training strategies to reduce computational requirements while maintaining generation quality. 2) integration of physical priors and domain knowledge to improve the realism of generated videos. 3) exploration of hybrid approaches that combine the strengths of diffusion models with other generative frameworks. 4) investigation of scalable solutions for real-world applications, particularly in domains such as virtual reality and autonomous systems.

### v2

@@ 4.5 Long video
The world exists in three spatial dimensions plus time, while a video represents a 2D projection of this 3D world. Consequently, video diffusion models should adhere to certain 3D constraints. These models generate multiple frames simultaneously and capture inter-frame relationships through attention layers spanning different frames. Existing research, such as SORA (Li et al., 2024i), demonstrates that by training on large-scale datasets, networks can implicitly learn 3D priors embedded within the data.
@@ 4.6.1 Training on 3D dataset
SV3D (Blattmann et al., 2023a) fine-tunes a pre-trained stable video diffusion model on orbiting videos of 3D synthetic objects from the GSO (Downs et al., 2022) and OmniObject3D (Wu et al., 2023d) datasets. These orbiting videos include both static orbits, where the camera circles an object at regularly spaced azimuths with a fixed elevation, and dynamic orbits, where the camera trajectory is irregular. Similarly, V3D (Chen et al., 2024g) fine-tunes on the Objaverse dataset, leveraging 360-degree orbiting videos, along with proprietary datasets featuring similar setups (Junlin Han, 2024; Melas-Kyriazi et al., 2024).
@@
CineMaster (Wang et al., 2025a) proposes a customized data labeling pipeline that leverages existing depth prediction (Yang et al., 2024e; f), instance segmentation (Ravi et al., 2024) and scene reconstruction (Zhang et al., 2024d) models to generate labeled data. This enables a 3D-aware model to train on large-scale, in-the-wild video data.
@@ 4.6.2 Architecture for 3D diffusion models
Recently, multi-view diffusion models (Shi et al., 2023d; Tang et al., 2023; Yang et al., 2024b; Liu et al., 2024b; f; Hu et al., 2024a; Kant et al., 2024; Yang et al., 2024d; Oh et al., 2023; Tang et al., 2024) have gained significant attention for static 3D scene generation. By employing cross-frame attention, these models effectively enhance the spatial relationships necessary for 3D consistency. For static 3D scene generation, epipolar-constrained attention (Xu et al., 2024b; Huang et al., 2023c; Zheng et al., 2024a) further enforces strict adherence to 3D constraints, ensuring greater structural accuracy.
@@
In DiT-like structures (Gao*et al., 2024; Chou et al., 2024; Liang et al., 2024b; Gu et al., 2025; Wang et al., 2025a), cross-frame attention is inherently achieved as tokens for different temporal and spatial locations are jointly mixed through transformer, allowing the model to learn 3D priors directly from data.
@@ 4.6.3 Camera conditioning
  • MLP-based embedding. A straightforward method is to use a fully connected multi-layer perceptron (MLP) to embed camera parameters. These parameters are often represented as elevation and azimuth angles, as in SV3D (Voleti et al., 2024; Yang et al., 2024c), panning and zooming movement (Yang et al., 2024h), or as a relative camera extrinsic matrix (Van Hoorick et al., 2024; Wang et al., 2024k; 2025a). This embedding serves as a global condition, typically introduced into the diffusion model via cross-attention layers.
@@ 5.1 Video representation learning
As shown by previous works in image generation (Chi et al., 2024), a pretrained video generative model might be able to serve as prior, loss, or regularize in learning video representations. Recently, a pioneer work (Yeh et al., 2024a) employs image and video generative models for hateful visual content recognition. It queries well-trained SOTA image and video generative models to automatically scale up training data pairs for hateful visual content recognition, achieving superior performance.
@@ 5.2 Video retrieval
Video retrieval, mainly video-text retrieval (VTR), which involves cross-modal alignment and abstract understanding of temporal images (videos), has been a popular and fundamental task of language-grounding problems (Wang et al., 2020a; b; 2021; Yu et al., 2023b). Most existing conventional video-text retrieval frameworks (Yu et al., 2017; Dong et al., 2019; Zhu & Yang, 2020; Miech et al., 2020; Gabeur et al., 2020; Dzabraev et al., 2021; Croitoru et al., 2021; Liu et al., 2025b; Wang et al., 2025c) focus on learning powerful representations for video and text and extracting modality-dependent representations with modality-dependent encoders. Inspired by the success of self-supervised pretraining methods (Devlin et al., 2019; Radford et al., 2019; Brown et al., 2020) and vision-language pretraining (Li et al., 2020; Gan et al., 2020; Singh et al., 2022) on large-scale unlabeled cross-modal data, recent works (Lei et al., 2021; Cheng et al., 2021; Gao et al., 2021; Ma et al., 2022; Park et al., 2022; Wang et al., 2022a; b; Zhao et al., 2022; Gorti et al., 2022; Luo et al., 2022; Liu et al., 2022d; a; Cao et al., 2022) have attempted to pretrain or fine-tune video-text retrieval models (Gorti et al., 2022; Lei et al., 2021).
With the rapid progress in video generation, leveraging the advancement in video generation (Chen et al., 2024c; Muaz et al., 2023) to advance video retrieval by reranking (Zhou et al., 2024e; Dutta et al., 2021) or other methods becomes more and more promising. On the other side, retrieval augmented generation could also help faithful video generation (Arefeen et al., 2024; Rosa, 2024; Tevissen et al., 2024; He et al., 2023a).
@@ 5.4.1 Video diffusion for 3D generation
V3D (Chen et al., 2024g) leverages the predictions of fine-tuned video diffusion models as direct supervision to generate 3D Gaussian splatting or meshes. SV3D (Voleti et al., 2024) employs the outputs of a fine-tuned Stable Video Diffusion model to supervise a Neural Radiance Field (NeRF) in the coarse stage, then uses SDS loss to refine high-frequency details of DMTet (Shen et al., 2021) initialized from the coarse stage. IM-3D (Melas-Kyriazi et al., 2024) adopts an instructive NeRF-to-NeRF (Haque et al., 2023) pipeline for distillation. This involves iteratively adding noise to rendered results and using a diffusion model to denoise the images, which then serve as targets for updating the 3D representation. VFusion3D (Junlin Han, 2024) demonstrates the potential of fine-tuned video diffusion models as data generators for creating large-scale 3D datasets. These datasets can be used to train feed-forward 3D generation models. Hi3D (Yang et al., 2024c) follows a coarse-to-fine 3D generation pipeline. It first generates low-resolution orbiting videos using a text-to-video model, then refines these into high-resolution orbiting videos with a video-to-video model. Both models are fine-tuned from pre-trained video diffusion models with camera condition extensions. The high-resolution orbiting videos are subsequently used to reconstruct 3D meshes. ReconX (Liu et al., 2024a) fine-tunes video diffusion models for sparse-view novel view synthesis, enabling efficient reconstruction of 3D scenes from limited input views. With recent advancements in large reconstruction models (LRMs) (Hong et al., 2023b; Zhang et al., 2024e), Wonderland (Liang et al., 2024b) demonstrates that by training a feed-forward 3D reconstruction model from the latent space of a video generative model, remarkable scene generation results can be achieved. Some examples are shown in Fig. 25.
@@ 5.4.2 Video diffusion for 4D generation
Several approaches leverage video diffusion models to train dynamic Neural Radiance Fields (NeRFs) or other 4D representations: MAV3D (Singer et al., 2023), 4D-fy (Bahmani et al., 2024c), Dream-in-4D (Zheng et al., 2024b), Align-Your-Gaussians (Ling et al., 2024a), and Animate124 (Zhao et al., 2023e) apply Score Distillation Sampling (SDS) loss along the temporal axis to train dynamic NeRFs using video diffusion models. These methods enable the models to learn motion dynamics directly from video diffusion. Vidu4D (Wang et al., 2024g) and STAG4D (Zeng et al., 2024b) propose robust reconstruction pipelines to reconstruct dynamic Gaussian Surfels or Gaussian splatting with monocular videos generated by video diffusion models. PLA4D (Miao et al., 2024a) animates static 3D Gaussian splatting generated by 3D diffusion models, using video diffusion predictions combined with a motion alignment model. TC4D (Bahmani et al., 2024a) uses SDS loss to supervise dynamic NeRFs, incorporating a trajectory-conditioned video model that disentangles global and local motion components, facilitating the generation of large-scale motions.
Dynamic motion knowledge can also be combined with spatial 3D knowledge from multi-view diffusion models: EG4D (Sun et al., 2024a) use a combination of SVD and SV3D models to directly generate multi-view videos, which are then used to supervise a 4D Gaussian splatting. 4Real (Yu et al., 2024a) generates 4D videos by using video diffusion model to supervise fixed-viewpoint video and freeze-time video with SDS loss. The use of deformable 3D GS ensures the temporal and view consistency. DreamGaussian4D (Ren et al., 2023) employs a training-free video-to-video pipeline that uses pre-trained video diffusion models to refine the dynamic textures of generated mesh sequences.
@@
6 Ethical Considerations
While video generation has been largely advanced by high-quality datasets (LAION, 2023; Miech et al., 2019; Yang et al., 2024a; Chen et al., 2024d), advanced model architectures (Lee et al., 2022; Seo et al., 2022), and training techniques (Wan et al., 2025; teamMeta, 2024), they (Wan et al., 2025; Zheng et al., 2024c; teamMeta, 2024) simultaneously raise critical ethical concerns that demand careful consideration.
Deepfakes and Misinformation. Creating realistic fake videos of real people without consent raises serious concerns about non-consensual intimate imagery (Zhao et al., 2021; Wu et al., 2022c; Rössler et al., 2018; Liu et al., 2025a), political manipulation, and spreading false information. This technology can undermine trust in authentic media and enable harassment campaigns. While image-based deepfake detection (Dang et al., 2020) has been successful, video deepfake video detection (Kundu et al., ) remains a serious challenge to this area.
Consent and Privacy. Using someone’s likeness, voice, or personal content to generate videos without explicit permission violates their privacy rights and autonomy (Rosenblat et al., 2025; Zhang et al., 2024f; Chen et al., 2024a). This extends to using training data that may contain copyrighted or private material without proper authorization.
Bias and Representation. AI models trained on biased datasets can perpetuate harmful stereotypes (Gallegos et al., 2024; Chen et al., 2024a), underrepresent certain groups, or generate content that reinforces discriminatory patterns. The bias in generated videos (Nadeem et al., 2025) could amplify existing social inequalities and marginalize vulnerable communities.
Legal and Regulatory Challenges. Current laws (Goanta et al., 2023) struggle to address AI-generated content, creating uncertainty around liability, intellectual property rights (Zhang et al., 2024f; Chen et al., 2024a), and regulatory compliance. This legal ambiguity can enable harmful uses while hindering beneficial applications.
Transparency and Disclosure. There’s an ongoing debate about when and how AI-generated content should be labeled. Clear disclosure (Chang et al., 2024b; Yoo et al., 2024; Chen et al., 2024a) helps viewers make informed decisions about the content they consume, but enforcement remains challenging. While applying watermarks on generated video (Hu et al., 2024b) could be a potential solution, more and more approaches should be explored to emphasize these considerations.
Quality Control and Safety. Video generation systems may produce harmful, offensive, or dangerous content if not properly constrained. Ensuring robust content filtering (Poppi et al., 2024; Yoon et al., 2025; Liang et al., 2025b; Liu et al., 2025a) and safety measures is crucial to prevent the creation of videos that could cause harm or violate policies.
Computational Resources and Environmental Impact. Training and running video generation models requires significant computational power, leading to substantial energy consumption and environmental costs (Nag et al., 2024) that should be considered in the development and deployment of these technologies.
Despite that, researchers have also proposed benchmark datasets (Miao et al., 2024b; Pang et al., 2024; Chen et al., 2025b; Yeh et al., 2024b; Wang et al., 2025b) to access the safety of text-to-video generative models.
@@ 7 Conclusion and outlook
Corresponding to these challenges, we anticipate several promising research directions:
  • Efficient Architectures and Training Strategies. The computational demands of video generation models present a significant barrier to widespread adoption and real-time applications (Bolya & Hoffman, 2023; Li et al., 2024j; Kim et al., 2023b; Sun et al., 2025; Kahatapitiya et al., 2024). Future research should focus on developing token-level optimizations such as adaptive token selection and learned pruning mechanisms that can identify and eliminate redundant spatial-temporal information while preserving critical motion dynamics. Additionally, hierarchical generation frameworks that decompose video synthesis into coarse-to-fine stages can leverage temporal coherence to reduce computational overhead. Novel training paradigms including progressive training (Wan et al., 2025), curriculum learning for temporal sequences, and specialized attention mechanisms for video data offer promising avenues for achieving better quality-efficiency trade-offs.
  • Integration of Physical Priors and Domain Knowledge. Current video generation models often produce visually plausible but physically implausible results (Kang et al., 2025a), limiting their applicability in domains requiring realistic motion and object interactions. Future research should systematically incorporate physical understanding through physics-informed architectures that explicitly model laws such as conservation of momentum, gravity, and collision dynamics (Bansal et al., 2024; Kang et al., 2025b; Bansal et al., 2025). For specialized applications such as medical video generation (Cao et al., 2024; Li et al., 2024f) or autonomous driving simulations (Fu et al., 2024b), incorporating domain-specific knowledge through structured priors can enhance both realism and utility. This requires developing modular frameworks that can easily integrate external knowledge sources and differentiable physics simulators into the generation pipeline.
  • Hybrid Generative Approaches. The integration of diffusion models with other generative frameworks (Li et al., 2025b; Xie et al., 2025) presents opportunities to leverage the complementary strengths of different approaches while mitigating their individual limitations. Future research should explore combinations of diffusion models with GANs (Wang et al., 2023j) for improved training stability and generation speed, autoregressive models (Lee et al., 2022) for better temporal coherence, and variational autoencoders (Pandey et al., 2022) for enhanced controllability. Additionally, investigating unified frameworks (Guo et al., 2025b; Zhang et al., 2025) that can seamlessly switch between different generative paradigms based on content requirements and computational constraints offers promising directions for more versatile video generation systems.
  • Model Safety and Ethical Considerations. As video generation technology becomes more sophisticated and accessible, ensuring model safety and addressing ethical concerns (Miao et al., 2024b; Pang et al., 2024; Chen et al., 2025b; Yeh et al., 2024b; Wang et al., 2025b; Hu et al., 2024b) becomes increasingly critical, such as copyrights, data consent, privacy, and biases.